OpenStudy (anonymous):

If 125 cal of heat is applied to a 60.0 g piece of copper at 22.0 Celsius, what will the final temperature be? The specific heat of copper is 0.0920 cal/g

OpenStudy (anonymous):

\[q=mc \Delta T\] plug in numbers and solve for Delta T, then you want to add that to the temp of the copper

OpenStudy (anonymous):

HUH?? I have no idea what you are talking about. Care to break it down for an old lady?? :)

OpenStudy (anonymous):

\[125Cal=(60.0g )(0.0920 \frac{Cal}{g}) \Delta T\] Now you have to solve for delta T \[\frac{125 Cal}{(60.0 g)(0.0920 \frac{Cal}{g}}= \Delta T\] \[\Delta T= 22.64\] Now that you know what delta t equals you must plug it into another equation \[\Delta T=T_f-T_i\] we plug in the rest of our known values \[22.64=T_f-22\] solve for T_f \[44.64 C =T_f\]

OpenStudy (anonymous):

Thank you. Thank you!

OpenStudy (anonymous):

No problem good luck to you!